# RiTeK: A Dataset for Large Language Models Complex Reasoning over Textual Knowledge Graphs

### v1

@@ Abstract
Answering complex real-world questions often requires accurate retrieval from textual knowledge graphs (TKGs). The scarcity of annotated data, along with intricate topological structures, makes this task particularly challenging. As the nature of relational path information could enhance the inference ability of Large Language Models (LLMs), efficiently retrieving more complex relational path information from TKGs presents another key challenge. To tackle these challenges, we first develop a Dataset for LLMs Complex Reasoning over Textual Knowledge Graphs (RiTeK) with a broad topological structure coverage. We synthesize realistic user queries that integrate diverse topological structures, relational information, and complex textual descriptions. We conduct rigorous expert evaluation to validate the quality of our synthesized queries. And then, we introduce an enhanced Monte Carlo Tree Search (MCTS) method, Relational MCTS, to automatically extract relational path information from textual graphs for specific queries. Our dataset mainly covers the medical domain as the relation types and entity are complex and publicly available. Experimental results indicate that RiTeK poses significant challenges for current retrieval and LLM systems, while the proposed Relational MCTS method enhances LLM inference ability and achieves state-of-the-art performance on RiTeK.
RiTeK: A Dataset for Large Language Models Complex Reasoning over Textual Knowledge Graphs
Jiatan Huang1thanks: indicates equal contribution, Mingchen Li211footnotemark: 1, Zonghai Yao2, Zhichao Yang2, Yongkang Xiao3 Feiyun Ouyang1, Xiaohan Li4 , Shuo Han1, Hong Yu1,2,5 1University of Massachusetts, Lowell 2University of Massachusetts, Amherst 3University of Minnesota 4Rollins School of Public Health, Emory University 5UMass Chan Medical School 1{Jiatan_Huang, Feiyun_Ouyang, shuo_han,Hong_Yu}@uml.edu 2{mingchenli, zonghaiyao, zhichaoyang}@umass.edu 3xiao0290@umn.edu,4xiaohan.li@emory.edu
@@ 1 Introduction
Although large language models (LLMs) have made a significant stride in natural language processing (NLP), complex question answering remains a challenge. Medical professionals, for instance, often need to express complex information that combines flexible inputs with specific, structured constraints. Consider the query: “Which organ or tissue function, that circulates mother and fetus blood, is affected by Fetal Distress?” compared to the simpler version: “What does Fetal Distress affect?” Accurately addressing such complex queries is critical, as it directly impacts healthcare diagnoses and treatment plans.
Figure 1: An example demonstrating the input, expected outputs and system outputs (from GPT-4, GPT-COT, Random Walk, MCTS, Relational MCTS (RMCTS)).
@@
To effectively answer these queries, organizing the underlying knowledge using textual knowledge graphs (TKGs) becomes essential. TKGs integrate unstructured data—such as textual descriptions of nodes (e.g., the definition of the medical term Placental Circulation)—with structured data, like the relationships between entities within the graph (e.g., the relationship between Fetal Distress and Placental Circulation is affects). This integration enables TKGs to represent comprehensive knowledge tailored to specific applications, rendering them invaluable, especially in the medical field, where accuracy and reliability are critically important.
@@
However, existing datasets Wu et al. (2024b, a) exhibit one or more of the following limitations: they are overly simplistic, involving only 1-2 hop reasoning paths; they lack diverse topological structure templates  [^1] A topological structure is a graph that is an abstract of the query graphs of the same pattern, as shown in Li and Ji (2022). and rich relation types; or they fail to incorporate complex constraints  [^2] Constraints are particularly important in KBQA as they help filter out irrelevant information from large knowledge bases, narrowing the search space and improving both efficiency and accuracy. Consequently, these datasets fall short in addressing the complexity of retrieval tasks involving TKGs, especially within the medical domain, where queries require more complex multi-hop reasoning, diverse topological structure templates, and multiple interdependent constraints. Moreover, the absence of textual properties in existing TKGs limits their effectiveness in delivering comprehensive answers.
In addition to the limitations of the TKG datasets themselves, another major challenge lies in how to leverage the useful information embedded in such complex TKGs to enhance the ability of LLMs to answer questions that involve complex reasoning paths within TKGs. As shown in Figure 1, both GPT-4333https://openai.com/index/gpt-4/ and Chain-of-Thought (COT) Wei et al. (2022) fail to generate the correct answer, Nitric oxide. Recent work suggests that utilizing path information within complex KGs can significantly enhance LLM reasoning abilities Li et al. (2024) and provide interpretability. Therefore, exploring accurate and efficient path retrieval with complex topological structures in TKGs is another key focus of the current research community.
@@
To address these gaps, we first present a large-scale complex reasoning dataset over textual knowledge graphs (RiTeK) within the medical domain. In this progress, one primary technical challenge we address is the accurate simulation of user queries with different reasoning types (e.g., six topological structures in Figure 1) within TKGs, ensuring that these queries are relevant and reflective of real-world medical scenarios involving patients, doctors, and medical scientists. This challenge arises from the interdependence of textual and relational information, the complexity of medical terminology and relationships, and the lack of textual descriptions of medical terms. We refer to the framework of Wu et al. (2024b) to simulate user queries and construct precise ground truth answers. However, our focus is primarily on the medical domain, incorporating richer topological structures that extend beyond the traditional 2- and 3-hop structures to better reflect real-world scenarios in the medical domain. Additionally, the textual descriptions of each node are more detailed, enhancing the overall context and understanding. With RiTeK, we delve deeper into retrieval tasks on TKGs, evaluate the capability of current retrieval systems, and provide insights for future advancement.
To address the second challenge and enhance the inference capabilities of LLMs in TKGs within the medical domain, we adapt the Monte Carlo Tree Search Chaslot (2010) to TKGs, and propose an improved Monte Carlo Tree Search, called Relational MCTS to dynamically retrieve relevant relational information for input queries. Our experiments demonstrate that relational MCTS is effective to retrieve the relational path information and enhance the traditional LLM, such as GPT-4.
Our main contributions are summarized as follows. (1) We propose RiTeK, a new QA dataset with complex reasoning over textual knowledge graph in the medical domain. (2) We propose a novel relational MCTS to retrieve the relational information and improve the inference abilities of LLMs within medical TKGs. (3) We delve deeper into retrieval tasks on RiTeK, evaluate the capability of current retrieval systems, and provide insights for future advancement. We also conduct a thorough analysis of our proposed method, including an ablation study, demonstrating the robustness of our framework.
@@ 2 Related Work
Datasets of Question Answering over Document. This area of research centers on extracting answers from document sources Rajpurkar (2016); Dunn et al. (2017); Joshi et al. (2017); Trischler et al. (2016); Welbl et al. (2018); Yang et al. (2018); Jin et al. (2021, 2019); Hendrycks et al. (2020). For example, SQuAD Rajpurkar (2016) assesses a model’s ability to interpret and retrieve answers from a single document, focusing on comprehension within a defined context. PubMedQA Jin et al. (2019) focuses on understanding and reasoning within the context of complex biomedical scientific texts. MedQA-CS Yao et al. (2024) aims at simulating authentic clinical scenarios encountered in the clinical examination tasks of medical education. However, unstructured QA datasets often fall short in providing the depth needed for complex relational reasoning to effectively tackle complex user inquiries. In contrast, our research involves queries that demand more complex relational reasoning, challenging the model’s ability to navigate and utilize structured information effectively.
Datasets of Question Answering over Knowledge Graph. The structure QA dataset challenges models to retrieve answers from the structured database, such as knowledge graph Zhang et al. (2018); Yih et al. (2016); Gu et al. (2021); Bao et al. (2016); Trivedi et al. (2017). For example, MetaQA Zhang et al. (2018) challenges models to generate the relational path with multi-hops. To test the models’ abilities to decompose the constraint information in the queries, WebQuestionsSP Yih et al. (2016) is proposed. GrailQA Gu et al. (2021) aims to facilitate the answering of more complex questions, as it allows queries to involve up to four relations and optionally includes functions such as counting, superlatives, and comparatives. However, these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information.
Datasets of Question Answering over Textural Knowledge Graph. To integrate textual information into knowledge graphs and queries, STARK (Prime, Amazon, Mag) Wu et al. (2024b) is proposed. To the best of our knowledge, Stark is the only work that focuses on combining relational information with textual information in the question answering over TKGs. However, this dataset lacks sufficient topological structure coverage, hindering the ability to handle complex queries, particularly in the medical domain. The absence of detailed node descriptions further challenges the model’s ability to understand the query information. RiTeK addresses these issues by incorporating richer topological structures and more extensive textual information into knowledge graphs and queries, resulting in more comprehensive and nuanced responses with deeper insights drawn from abundant textual data.
@@ 3 Problem Statement
Textural Knowledge Graph
A Textural Knowledge Graph (TKG) is defined as a graph $\mathcal{G}=(\mathcal{R},\mathcal{E},\mathcal{D})$, where $\mathcal{E}$ denotes a set of entities and $\mathcal{R}$ denotes the set of relations among these entities. In a TKG, the entities and relations are usually organized as facts and each fact is defined as a triplet $(h,r,t)$ where $h,t\in\mathcal{E}$ and $r\in\mathcal{R}$ denote the head entity, tail entity and the relation between the two entities, respectively. Each entity $e$ ($e=h$ or $e=t$) in $\mathcal{G}$ has a textual document $d^{e}\in\mathcal{D}$ describing the entity information.
Complex Question Answering over Textural Knowledge Graph
Given a textual knowledge graph $\mathcal{G}$ and input query $q$, the model is expected to generate the answers $a\in\mathcal{E}$, which satisfy the relational constraints defined by the structure of $\mathcal{G}$ as specified in $q$, and the associated document $d^{e}$ need to satisfy the textual requirements in $q$.
Textural Triple Graph
Unlike traditional knowledge graphs, where each node represents an entity and each edge denotes the relationship between nodes, in the textual triple graph, each node corresponds to a triple (head entity, relation, tail entity) along with the textual description of each entity. In this context, the relation indicates whether the two triples are connected. To be specific, let $\mathcal{G*}=(V,E)$ denote a graph consisting of a set of node $V$ and a set of edges $E\in V\times V$. We denote by $n$ the number of nodes in $\mathcal{G}$ and by $m$ its number of edges. Each node $v=(h,r,t,T(h),T(t))\in V$, $T(*)$ is the textual description of entity.
@@ 4 Dataset for LLMs Complex Reasoning over TKGs (RiTeK)
Table 1: Datasets Statistics of constructed medical textual knowledge graphs. # Coverage refers to textual description coverage of each node. # Entities is the number of entity types. # Relation is the number of relation types. As the textual information for the provided nodes is difficult to analyze statistically, we have not provided the statistical information for Stark Mag and Amazon.
@@ 4.1 Medical Textual Knowledge Graph Construction
We construct two medical TKGs based on PharmKG Zheng et al. (2021) and ADint Xiao et al. (2024), as the increased number of entity and relation types introduces significant challenges for path retrieval in the question answering over textual knowledge graph. To enhance the entity attributes, we incorporate textual details from various databases, including Ensembl, UMLS, and Mondo Disease Ontology. As shown in Table 1, our constructed TKGs provide greater node textual coverage, along with more entity types and relation types. For further details on these two medical TKGs, please refer to Appendix A.2.
@@ 4.2.1 Overview
We developed two question answering datasets RiTeK PharmKG and RiTeK ADint based on two textual knowledge graphs for complex reasoning. These datasets notably feature queries that integrate relational and textual knowledge, incorporating relational templates with broader coverage and higher instance rates. Additionally, to enhance their applicability in practical scenarios, these queries mimic real-life query patterns, exhibiting a natural-sounding quality and flexible formats. Specifically, RiTeK-PharmKG consisting of 10,235 synthesized queries. To maximize the coverage of the question topology, we generate the queries following the six types of topological structure (e.g., multi-hop and multi-hop with constraints). For the synthesized queries, we developed 68 relational templates, crafted by a medical expert and detailed in Appendix C.1, to encompass various relation types and ensure practical relevance. The instance rate of 11.33, which is higher than that of the current TKG dataset Start (Amazon, Mag, and Prime), highlights the higher diversity of this dataset. RiTeK-ADint consisting of 5322 synthesized queries also covering 6 topological structures, with 58 relational templates, and detailed in Appendix C.2. To capture diverse language styles used by different users, same as Stark, we also simulate three distinct roles: medical scientist, doctor, and patient. We divide the synthesized queries on each dataset into training, validation, and testing subsets, with the ratios detailed in Table  2, we use random splits to divide the data. Further details on the scale of our QA benchmarks can be found in Table 2.
4.2.2 Construction Pipline
@@
We present the pipeline that generates the large-scale medical QA datasets on TKGs. The core idea is to entangle relational information and textual properties in to the query, and accurately construct ground truth answers with more topological structures. The construction of the QA datasets (Figure 2) generally involves five steps, and the specific processes vary depending on the characteristics of each dataset. These steps are as follows:
@@
Relational Template Construction. As shown in Figure 2 Step1. We first created templates based on the 6 designed topological structures Li and Ji (2022), and these templates were evaluated by medical experts to ensure their practical relevance and value. Afterward, the topological structures are instantiated conceptually with experts. For instance, for the topological structure Head entity–relation–tail entity, the "(antibiotic) causes <pathologic function>" is a valid and common medical relation template, as antibiotics, particularly penicillin and cephalosporins, are well-known for triggering drug hypersensitivity reactions. This makes it a medically reasonable and frequently observed relationship. We then converted these relation templates into specific relationship queries, such as "Anti-Bacterial Agents causes pathologic function." Since each query could correspond to one or more candidate entities, we matched the queries with the textual KG to obtain k candidate entities.
@@
Extracting Textual Properties. As shown in Figure 2 Step2. For the k candidate answers that meet the relationship criteria, we select one entity as the gold answer and use GPT-4 to extract textual properties from the entity’s associated document. For instance, in the relationship "Anti-Bacterial Agents causes pathologic function," we selected "Hypersensitivity" as the gold answer and extracted textual properties, as shown in Figure 2 (Texutal properties). These textual properties expand upon the concept of hypersensitivity, highlighting its key characteristics, which makes it more likely to fulfill the inquirer’s needs.
@@
Combining Textual and Relational Information. As shown in Figure 2 Step3. After obtaining the relationship templates and textual properties, we combine these components to synthesize the queries. We chose GPT-4 as the LLM for query synthesis, as it excels at generating natural, human-like questions. Additionally, we optimized the prompt and incorporated instructions for different personas to make the queries more diverse and realistic. This approach enhances the quality of our dataset and increases the demands on our model’s reasoning capabilities.
@@
Filtering Additional Answers. As shown in Figure 2 Step4. In addition to the gold answer from which the textual properties are extracted, we need to evaluate whether other candidates meet the requirements of the query in order to include them in the final answer set. We use multiple LLMs to assess whether each candidate’s description satisfies the textual requirements of the query. Only candidates that pass validation by all LLMs will be added to the final answer set.
@@
Human Evaluation. We invited four medical experts to evaluate 100 synthetic queries sampled from two datasets. The evaluation was conducted using a 5-point Likert-like scale across three dimensions. Naturalness refers to how grammatically and logically human-like the queries sound. Diversity assesses whether the queries exhibit complex logical structures and encompass multiple entities, relations, and textual requirements. Practicality measures the real-world applicability of the generated queries and their likelihood of being used in everyday scenarios. The scores were ultimately converted into percentages representing the rates of Positive and Acceptable responses. As shown in the Table 3, our two datasets received high ratings across all three dimensions. The average acceptable rates across the three dimensions were 95.6%, 99.3%, and 82.8%, respectively.
The experts observed that certain queries present issues related to practicality. For instance, the query, What chemical, associated with the metabolism and pharmacokinetics of the gene ’mif’, is also soluble in water? demonstrates a misalignment between the question and the answer. The proposed answer, nitric oxide, lacks sufficient evidence to directly support its involvement in the metabolism of the ’mif’ gene. Furthermore, queries such as What is a disease that is an ancestor of cerebrovascular disorders and can result in brain ischemia? are more suitable for medical education contexts rather than for everyday question-answering scenarios. In contrast, queries like What binding of TF to channels affects the expression/production of NAT1, resulting in the chemical that helps prevent megaloblastic anemia? provide sufficient relational context and textual clarity, making it easier to arrive at the correct answer.
@@
RiTeK-PharmKG 77.3/93.3 97.3/99.3 65.3/86.0
RiTeK-ADint 80.2/97.9 95.9/99.3 65.9/79.6

### v3

@@ Abstract
Answering complex real-world questions in the medical domain often requires accurate retrieval from medical Textual Knowledge Graphs (medical TKGs), as the relational path information from TKGs could enhance the inference ability of Large Language Models (LLMs). However, the main bottlenecks lie in the scarcity of existing medical TKGs, the limited expressiveness of their topological structures, and the lack of comprehensive evaluations of current retrievers for medical TKGs. To address these challenges, we first develop a Dataset [^1] The dataset is available here: https://github.com/ToneLi/Medical-Textual-KG-Reasoning-Benchmark/tree/main for LLMs Complex Reasoning over medical Textual Knowledge Graphs (RiTeK), covering a broad range of topological structures. Specifically, we synthesize realistic user queries integrating diverse topological structures, relational information, and complex textual descriptions. We conduct a rigorous medical expert evaluation process to assess and validate the quality of our synthesized queries. RiTeK also serves as a comprehensive benchmark dataset for evaluating the capabilities of retrieval systems built upon LLMs. By assessing 11 representative retrievers on this benchmark, we observe that existing methods struggle to perform well, revealing notable limitations in current LLM-driven retrieval approaches. These findings highlight the pressing need for more effective retrieval systems tailored for semi-structured data in the medical domain.
RiTeK: A Dataset for Large Language Models Complex Reasoning over Textual Knowledge Graphs in Medicine
Jiatan Huang1thanks: indicates equal contribution, Mingchen Li211footnotemark: 1, Zonghai Yao2, Dawei Li3, Yuxin Zhang2, Zhichao Yang7 , Yongkang Xiao5 Feiyun Ouyang8, Xiaohan Li6 , Shuo Han8, Hong Yu2,4,8 1University of Connecticut 2University of Massachusetts, Amherst 3School of Computing, and Augmented Intelligence, Arizona State University 4UMass Chan Medical School,5University of Minnesota 6Rollins School of Public Health, Emory University 7Optum AI 8University of Massachusetts, Lowell
@@ 1 Introduction
Although large language models (LLMs) have made significant strides in natural language processing (NLP), complex question answering still remains a challenge. Medical professionals, for instance, often need to express complex information that combines flexible inputs with specific, structured constraints. Consider the query, “Which organ or tissue function that circulates maternal and fetal blood is affected by Fetal Distress?” compared with the simpler version, “What does Fetal Distress affect?” Accurately addressing such complex queries is crucial, as it directly impacts healthcare diagnosis and treatment planning.
@@
To effectively answer these queries, organizing the underlying knowledge using medical TKGs becomes essential. TKGs integrate unstructured data, such as textual descriptions of nodes (e.g., the definition of the medical term Placental Circulation) with structured data, like the relationships between entities within the graph (e.g., the relationship between Fetal Distress and Placental Circulation is affects). This integration enables TKGs to represent comprehensive knowledge tailored to specific applications, rendering them invaluable, especially in the medical field, where accuracy and reliability are critically important.
@@
However, existing datasets Wu et al. (2024b, a) exhibit several critical limitations: they are overly simplistic, typically limited to 1-2 hop reasoning paths; they lack diverse topological structure templates [^2] The details of the topological structure are provided in the Appendix C. and rich relation types; or they fail to incorporate complex constraints [^3] Constraints are particularly important in KBQA as they help filter out irrelevant information from large knowledge bases, narrowing the search space and improving both efficiency and accuracy.. Consequently, these datasets are inadequate for capturing the complexity of retrieval tasks involving medical TKGs, where queries demand multi-hop reasoning, diverse topological structure templates, and multiple interdependent constraints. Moreover, the absence of textual properties in existing medical TKGs limits their effectiveness in delivering comprehensive answers.
@@
To bridge this gap, we introduce RiTek, a large-scale dataset for complex reasoning over medical TKGs. In this progress, one primary technical challenge we address is the accurate simulation of user queries with different reasoning types (e.g., six topological structures in Figure 1 within medical TKGs, ensuring that these queries are relevant and reflective of real-world medical scenarios involving patients, doctors, and medical scientists. This challenge stems from the interdependence between textual and relational information, the inherent complexity of medical terminology and relationships, and the limited availability of textual descriptions for medical terms. We refer to the framework of Wu et al. (2024b) to simulate user queries and construct precise ground-truth answers. We incorporate richer topological structures that extend beyond the traditional 2- and 3-hop reasoning patterns to better reflect real-world medical scenarios. Compared with datasets like BioKGBench Lin et al. (2024) and PrimeKGQA Yan et al. (2024), the queries in RiTek not only contains structural information that requires reasoning but also includes textual information related to the ground-truth answer, making the task more challenging. In addition, we enrich the textual descriptions of each node and incorporate more ontology structure.
The key features of RiTeK are summarized as follows: (1) it integrates rich ontological structures and comprehensive textual descriptions, with content quality rigorously validated by medical experts to ensure reliability; (2) it constructs queries that capture complex relational dependencies and nuanced linguistic variations; and (3) it introduces context-sensitive reasoning, where effective retrieval depends not only on a model’s reasoning ability but also on its semantic alignment with entity constraints embedded in the query.
Moreover, we systematically investigate the performance of existing retrieval systems on RiTeK and provide insights to guide future research. In particular, we identify key challenges in processing textual and relational data with complex ontology structures and in mitigating latency issues on large-scale SKBs containing millions of entities and relations.
@@ 2 Related Work
Datasets of Question Answering over Document. This area of research centers on extracting answers from document sources Rajpurkar (2016); Dunn et al. (2017); Joshi et al. (2017); Trischler et al. (2016); Welbl et al. (2018); Yang et al. (2018); Jin et al. (2021, 2019); Hendrycks et al. (2020). For example, SQuAD Rajpurkar (2016) assesses a model’s ability to interpret and retrieve answers from a single document, focusing on comprehension within a defined context. PubMedQA Jin et al. (2019) targets reasoning over complex biomedical literature. MedQA-CS Yao et al. (2024b) aims to simulate authentic medical examination scenarios in clinical education. However, existing unstructured QA datasets often lack the depth required for relational reasoning and failed to address complex user inquiries. In contrast, our research involves queries that demand more complex relational reasoning, challenging the model’s ability to navigate and utilize structured information effectively.
Datasets of Question Answering over Knowledge Graph. Structured QA datasets challenge models to retrieve answers from knowledge graphs, which serve as structured databases for factual reasoning Zhang et al. (2018); Yih et al. (2016); Gu et al. (2021); Bao et al. (2016); Trivedi et al. (2017). For instance, MetaQA Zhang et al. (2018) requires models to infer multi-hop relational paths across entities. To test the models’ abilities to decompose the constraint information in the queries, WebQuestionsSP Yih et al. (2016) is proposed. GrailQA Gu et al. (2021) aims to facilitate the answering of more complex questions, as it allows queries to involve up to four relations and optionally includes functions such as counting, superlatives, and comparatives. However, these datasets primarily focus on relational information; the absence of textual context restricts query diversity and limits the semantic expressiveness of reasoning within predefined relationships and entities.
Datasets of Question Answering over Textual Knowledge Graph. To integrate textual information into knowledge graphs and queries, the STaRK dataset (Prime, Amazon, Mag) Wu et al. (2024b) was proposed. To the best of our knowledge, STaRK remains the only dataset that integrates relational and textual information for question answering over TKGs. However, this dataset exhibits limited topological structure coverage, which restricts its ability to handle complex multi-hop queries, particularly in the medical domain. Furthermore, the lack of detailed node descriptions further impairs a model’s ability to comprehend query semantics. RiTeK addresses these limitations by incorporating richer topological structures and more extensive textual information into both knowledge graphs and queries. This integration leads to more comprehensive and nuanced responses, providing deeper insights drawn from abundant textual data.
@@ 3 Problem Statement
Textual Knowledge Graph
A Textual Knowledge Graph (TKG) is defined as a graph $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{D})$, where $\mathcal{E}$ denotes a set of entities and $\mathcal{R}$ denotes a set of relations among these entities. In a TKG, the entities and relations are usually organized as facts, and each fact is defined as a triplet $(h,r,t)$, where $h,t\in\mathcal{E}$ and $r\in\mathcal{R}$ denote the head entity, tail entity and the relation between the two entities, respectively. Each entity $e$ ($e=h$ or $e=t$) in $\mathcal{G}$ is associated with a textual document $d^{e}\in\mathcal{D}$ describing its information.
Complex Question Answering over Textual Knowledge Graph
Given a textual knowledge graph $\mathcal{G}$ and input query $q$, the model is expected to generate the answers $a\in\mathcal{E}$, which satisfy the relational constraints defined by the structure of $\mathcal{G}$ as specified in $q$, and the associated document $d^{e}$ needs to satisfy the knowledge required to solve $q$.
Textual Triple Graph
Unlike traditional knowledge graphs, where each node represents an entity and each edge denotes the relationship between nodes, in the textual triple graph, each node corresponds to a triple (head entity, relation, tail entity) along with the textual description of each entity. In this context, the relation indicates whether the two triples are connected. To be specific, let $\mathcal{G*}=(V,E)$ denote a graph consisting of a set of nodes $V$ and a set of edges $E\in V\times V$. We denote by $n$ the number of nodes in $\mathcal{G}$ and by $m$ its number of edges. Each node $v=(h,r,t,T(h),T(t))\in V$, where T()T(*) denotes the textual description of an entity.
@@ 4 Dataset for LLMs Complex Reasoning over TKGs (RiTeK)
Table 1: Dataset Statistics of constructed medical textual knowledge graphs. # Coverage refers to the proportion of nodes with textual descriptions. # Entities denotes the number of entity types, and # Relations indicates the number of relation types. As the textual information of the provided nodes is difficult to quantify statistically, we do not include the corresponding statistics for STaRK-Mag and STaRK-Amazon.
@@ 4.1 Medical Textual Knowledge Graph Construction
We construct two medical TKGs based on PharmKG Zheng et al. (2021) and ADint Xiao et al. (2024), as the increased number of entity and relation types introduces significant challenges for path retrieval in question answering over textual knowledge graphs. To enrich entity representations, we incorporate textual details from various databases, including Ensembl, UMLS, and Mondo Disease Ontology. As shown in Table 1, our constructed TKGs provide greater node textual coverage, as well as a larger variety of entity and relation types. For further details on these two medical TKGs, please refer to Appendix A.4.
@@ 4.2.1 Overview
We developed two question-answering datasets, RiTeK-PharmKG and RiTeK-ADint, based on textual knowledge graphs for complex reasoning. These datasets notably feature queries that integrate relational and textual knowledge, incorporating relational templates with broader coverage and higher instance rates. Additionally, to enhance their applicability in practical scenarios, these queries mimic real-world query patterns, exhibiting a natural-sounding quality and flexible formats. Specifically, RiTeK-PharmKG consists of 10,235 synthesized queries. To maximize the coverage of different question topologies, we generate the queries following the six types of topological structure (e.g., multi-hop and constrained multi-hop). For the synthesized queries, we developed 68 relational templates, crafted by medical experts and detailed in Appendix B.1, to encompass various relation types and ensure practical relevance. The instance rate of 11.33, which is higher than that of the current TKG dataset STaRK (Amazon, Mag, and Prime), highlights the greater diversity of our dataset. RiTeK-ADint consists of 5322 synthesized queries and covers 6 topological structures, with 58 relational templates. Further details are provided in Appendix B.2. To capture the diverse language styles used by different users, we follow STaRK and simulate three distinct roles: medical scientist, doctor, and patient. We divide the synthesized queries on each dataset into training, validation, and test subsets, with the ratios detailed in Table  2. Further details on the scale of our QA benchmarks can be found in Table 2.
4.2.2 Construction Pipeline
@@
We present the pipeline used to generates large-scale medical QA datasets on TKGs. The core idea is to intertwine relational information and textual properties within the queries, accurately constructing ground-truth answers that exhibit more complex topological structures. The construction of the QA datasets (Figure 1) generally involves five steps, and the specific processes vary depending on the characteristics of each dataset. These steps are as follows.
@@
Relational Template Construction. As shown in Figure 1 Step 1, we first created templates based on the 6 designed topological structures Li and Ji (2022), which were evaluated by medical experts to ensure their practical relevance and value. Afterward, the topological structures are instantiated conceptually with experts. For instance, for the topological structure Head entity–relation–tail entity, the "(antibiotic) causes <pathologic function>" is a valid and common medical relation template, as antibiotics, particularly penicillin and cephalosporins, are well-known for triggering drug hypersensitivity reactions. This makes it a medically reasonable and frequently observed relationship. We then converted these relation templates into specific relationship queries, such as "Anti-Bacterial Agents causes pathologic function." Since each query could correspond to one or more candidate entities, we matched the queries with the textual KG to obtain k candidate entities.
@@
Extracting Textual Properties. As shown in Figure 1 Step 2, for the k candidate answers that meet the relationship criteria, we select one entity as the gold answer and use GPT-4 to extract textual properties from the entity’s associated document. For instance, in the relationship "Anti-Bacterial Agents causes pathologic function," we selected "Hypersensitivity" as the gold answer and extracted its textual properties. These textual properties elaborate on the concept of hypersensitivity, highlighting its key characteristics, which make it more likely to meet the inquirer’s needs.
@@
Combining Textual and Relational Information. As shown in Figure 1 Step 3, after obtaining the relational templates and textual properties, we combine these components to synthesize the queries. We chose GPT-4 as the LLM for query synthesis, as it excels at generating natural, human-like questions. Additionally, we optimized the prompt and incorporated instructions for different personas to make the queries more diverse and realistic. This approach enhances the quality of our dataset and increases the demands on our model’s reasoning capabilities. For details on using GPT-4 to generate this query, please refer to Appendix A.5.
@@
Filtering Additional Answers. As shown in Figure 1 Step 4, in addition to the gold answer from which the textual properties are extracted, we need to evaluate whether other candidates meet the requirements of the query in order to include them in the final answer set. We use multiple LLMs to assess whether each candidate’s description satisfies the textual requirements of the query. Only candidates that pass validation by all LLMs will be added to the final answer set.
@@
Human Evaluation. We invited four medical experts to evaluate 1,000 synthetic queries sampled from two datasets. The evaluation was conducted using a 5-point Likert-like scale across three dimensions. Naturalness measures how grammatically correct and human-like the queries sound. Diversity assesses whether the queries exhibit complex logical structures and encompass multiple entities, relations, and textual requirements. Practicality evaluates the real-world applicability of the generated queries and their likelihood of being encountered in real clinical or everyday scenarios.
The scores were ultimately converted into percentages representing the rates of Positive and Acceptable responses. We found that the evaluation results provided by GPT-4 for our generated dataset were largely consistent with assessments from medical experts. For shorter queries, such as “What gene is inhibited by naloxone?”, GPT-4 noted the limited relational and textual information contained within and consequently assigned a lower Diversity score. Both GPT-4 and medical experts agreed that certain rare relationship types, such as “an ancestor of”, are infrequently encountered in everyday Q&A scenarios and are more common in medical education contexts. Only a very small number of queries exhibited issues with insufficient Practicality. The results of this evaluation are summarized in Table 3. The data in the table represents the Positive/ Acceptable rates (%) from GPT-4.
@@
RiTeK-PharmKG 81.80/99.60 81.6/99.40 67.4/97.8
RiTeK-ADint 81.20/99.20 74.80/100 68.60/96.60